### v3

# LINEAR PARAMETRIC MODEL CHECKS FOR FUNCTIONAL TIME SERIES

Abstract
The presented methodology for testing the goodness–of–fit of an Autoregressive Hilbertian model (ARH(1) model) provides an infinite–dimensional formulation of the approach proposed in Koul and Stute [41], based on empirical process marked by residuals. Applying a central and functional central limit result for Hilbert–valued martingale difference sequences, the asymptotic behavior of the formulated –valued empirical process, also indexed by is obtained under the null hypothesis. The limiting process is –valued generalized (i.e., indexed by ) Wiener process, leading to an asymptotically distribution free test. Consistency of the test is also proved. The case of misspecified autocorrelation operator of the ARH(1) process is addressed. The asymptotic equivalence in probability, uniformly in the norm of of the empirical processes formulated under known and unknown autocorrelation operator is obtained. Beyond the Euclidean setting, this approach allows to implement goodness of fit testing in the context of manifold and spherical functional autoregressive processes.
González–Manteiga, W., Ruiz–Medina. M.D., Febrero–Bande, M.
Key words. Generalized functional empirical processes, generalized Hilbert–valued Wiener process, manifold functional autoregressive processes, misspecification, strong consistency.
1 Introduction
Weakly dependent functional time series models have been extensively analyzed in the last few decades, supporting inference on stochastic processes, specially, in a state space framework (see, e.g., [1]; [22]; [28]; [33]; [36]; [38]). Confidence bands, kernel and parametric functional time series estimation lead to the analysis of outstanding problems like adaptive bandwidth selection, dimension reduction, change–point analysis, and functional principal component estimation, among others (see, e.g., [3]; [17]; [25]; [35]; [29]; [45]; [51]; [52]). Hypothesis testing is still a challenging topic in the functional time series literature. Testing independence constitutes a central subject in this topic (see, e.g., [2]; [23]; [31]; [49]). On the other hand, second–order properties like separability, stationarity, and linear correlation range are usually tested to specify the covariance operator family (see, for instance, [10]; [32]; [38]; [39]; [50]). Alternative contributions in testing in functional time series literature, related, for example, with normality and periodicity tests can be respectively found in [21] and [26]. Two–sample mean and change point tests have also been formulated in a Banach space context to improve shape analysis for random elements (see, e.g., [17]). See also the review on goodness–of–fit tests in González–Manteiga and Crujeiras [20].
Horváth et al. [31] formulate a test, based on the sum of the -norms of the empirical correlation functions, to check if a set of a functional observations are independent and identically distributed. Consistency is analyzed as well. To detecting changes in the autocorrelation operator in functional time series, Horváth and Reeder [33] derive a test based on change point analysis. In Horváth, Kokoszka and Rice [32], several procedures are considered to test the null hypothesis of stationarity in the context of functional time series. The properties of the tests under several alternatives are also studied. Approaches for testing the structural stability of temporally dependent functional observations are considered in Zhang et al. [51] . Kokoszka and Reimherr [38] propose a multistage testing procedure to determine the order of a functional autoregressive process. The test statistics involves the estimating kernel function in this linear model, and is approximately distributed as a chi–square with number of degrees of freedom determined by the number of functional principal components used to represent the data. Bootstrap methods constitutes a challenging topic in functional time series to testing trend-stationary, under the umbrella of Kwiatkowski–Phillips–Schmidt–Shin (KPSS) tests (see, e.g., Chen and Pun [8]). An interesting review on white noise tests in the context of functional time series is presented in Kim, Kokoszka and Rice [27], distinguishing between the formulation of such tests in the time or spectral domain, referring to the cases of strong white noise and general uncorrelated noise. The review also covers and extends several residual-based goodness-of-fit tests widely applied in functional data analysis. This survey provides, in addition, an interesting guide through simulations about implementation, benefits, and drawbacks of these tests.
In the more general context of functional regression
beyond the particular case of functional time series, restricting our attention to the separable Hilbert space framework, where in the case where and in the case special attention has been paid to significance testing (see, e.g., Hilgert, Mas and Andverzelen [24]; Chiou and Müller [9]; Bücher, Dette and Andwieczorek [5]). Some extensions to more complex hypothesis testing can be found in Delsol, Ferraty and Vieu [16]; Patilea, Sánchez–Sellero and Andsaumard [44]; García-Portugués, González-Manteiga and Febrero–Bande [19], among others. The last one adopts a marked empirical process based methodology, reducing the testing procedure to a multivariate problem by projection into the –sphere of A marked empirical process based approach is also adopted in Cuesta–Albertos et al. [13] in the context of goodness of fit tests for the functional linear model with scalar response and functional covariate. Specifically, the marked empirical process is indexed by random projections of the functional covariate. The weak convergence of the empirical process is obtained conditionally on a random direction. The almost surely equivalence of the test based on the original and projected covariate is proved. Calibration by wild bootstrap resampling is also achieved.
The present paper extends the formulation in Cuesta–Albertos et al. [13] to the context of autoregressive functional time series, i.e., functional linear model with functional response, and time–dependent functional covariates, in the case of error term given by a Hilbertian Strong White Noise. Specifically, an infinite–dimensional formulation of the involved marked empirical process indexed by a functional covariate is adopted. The limit behavior of the marginals of this process under the null hypothesis is obtained by applying a Central Limit result for Hilbert–valued martingale difference sequences (see Theorem 2.16 in Bosq [4]). A functional central limit result is also derived by applying a special case of an invariance principle based on Robbins-Monro procedure (see Theorem 2 in Walk [47]). A more general formulation can be obtained from the conditional central limit theorem in Hilbert spaces derived in Dedecker and Merlevede [15].
The functional central limit result formulated here characterizes the limiting process of the functional marked empirical process as a generalized Gaussian process taking values in a separable Hilbert space whose index set is also This process is identified in law with two–parameter Hilbert–valued generalized Brownian Sheet. The proposed goodness–of–fit test is based on one of its marginals indexed by the elements of The application of Theorem 4.1 in Cuesta–Albertos, Fraiman and Ransford [12] then allows to determine the corresponding asymptotic critical values from the probability distribution of a random projection of an infinite–dimensional Gaussian measure in with covariance operator given by the autocovariance operator of the innovation process in the ARH(1) model. Consistency of the formulated test is proved. The case of misspecified autocorrelation operator of the ARH(1) process is also addressed. The marks of the empirical process are approximated in this case from the consistent projection estimation of the autocorrelation operator of the ARH(1) process. The uniform, in asymptotic equivalence in probability in the norm of of the totally specified and misspecified functional empirical processes is proved. Thus, the limiting process in both cases coincides. Hence, one can proceed in the same way in the testing procedure applied under misspecified autocorrelation operator model.
The outline of the paper is the following. Preliminaries are provided in Section 2. Section 3.1 derives a Central Limit result to characterize the asymptotic behavior of the marginals of the –valued empirical process. A Functional Central Limit Theorem for the considered Hilbert–valued marked empirical process indexed by infinite–dimensional covariates is obtained in Section 3.2, leading to the design of the goodness–of–fit test. Sufficient conditions for the consistency of the proposed goodness–of–fit test are given in Section 4. The issue of misspecified autocorrelation operator of the ARH(1) process is addressed in Section 5. Section 6 discusses the differences arising in our ARH(1) context, in relation to the previously derived central limit results, in the framework of the functional regression model with functional covariate, in the cases of scalar and functional response. Finally, ongoing simulation analysis is announced, currently developed to illustrate the finite sample performance of the Kolmogorov–Smirnov and Cramér–von Mises–type statistical tests constructed over the considered Hilbert–valued empirical process, marked by functional residuals and indexed by
2 Preliminaries
In the following we denote by the space of zero–mean –valued random variables on the basic probability space with for every Let be a zero–mean ARH(1) process on the basic probability space satisfying
and the conditions given in Chapter 3 in Bosq (2000) for the existence of a unique stationary solution. Hence, denotes the autocovariance operator of and its cross–covariance operator. The –valued innovation process is assumed to be –valued Strong White Noise (–SWN) (see Definition 3.1 in Bosq, 2000). That is, is a sequence of independent and identically distributed zero–mean –valued random variables with
and functional variance that coincides with the trace norm of the autocovariance operator of We restrict our attention to the ARH(1) process family satisfying the following condition.
Assumption A1. is a strictly stationary ARH(1) process with SWN innovation such that
Under model (1), keeping in mind Assumption A1, the following orthogonality condition
holds for with and respectively denoting the equality in distribution and the almost surely (a.s.) equality.
Let us consider a centered non–degenerated Gaussian probability measure on From Theorem 4.1 in Cuesta–Albertos, Fraiman and Ransford [12], under Assumption A1, if satisfies Carleman condition, a class in of Borel sets of defines a separating class of if it contains a set of positive –measure. Thus, Lemma 1(d) in Escanciano [18] can be reformulated in an infinite–dimensional framework by considering, for each and for a given orthonormal basis of
leading to the definition of the separating class of sets Hence, under Assumption A1, one can equivalently express the orthogonality condition (3) as
for a given orthonormal basis of
The following lemma establishes the –valued martingale difference property of the sequence
Lemma 1
For every the sequence
is an –valued martingale difference with respect to the filtration where for
Proof. For each and
in view of the definition of the innovation process as for every with being the orthogonal projector into (see equation (2.55), and pp. 72–73 in Bosq [4]).
Along the paper, from (1), the following identity will be applied several times (see equation (3.11) in Bosq, 2000), for every
We also consider the following assumption, that, in particular, will be applied in the next section in the computation of second–order moments of –valued martingale difference for every
Assumption A2. Assume that is independent of for all
Second–order properties of –valued martingale difference sequence
This section computes, under Assumption A1–A2, the second–order moments of –valued martingale difference for every In particular, for any and
Remark 1
The strongly integrability of the marginals of the –valued martingale difference for every also allows the computation of means and conditional means of the elements of this sequence from their weak counterparts (Section 1.3 in Bosq [4]).
For every the functional variance can then be computed under Assumptions A1–A2 as follows:
where denotes the trace norm of a nuclear operator on Here, denotes the infinite–dimensional marginal probability measure induced by and denotes the event for each Here, we also have applied that under Assumptions A1–A2,
Note that, from equation (3.13) in Bosq [4], for the linear predictor one can consider the identity
Under Assumptions A1–A2, for each we compute the autocovariance operator of the sequence keeping in mind equation (5), as follows:
where In (7), we have again applied the strictly stationarity condition of under Assumption A1, leading to, for every
In a similar way, the covariance operator can be computed for every
where the notation must be understood componentwise, by projections, as follows from equation (9). That is, the infinite–dimensional vector has components defined by the respective minimum values between the two projections of the two functions
3 Limit results
In this section, a Central and Functional Central Limit result will be derived for the generalized functional empirical process given by
We will apply Lemma 2 below to obtain Theorem 1, providing the Gaussian limit distribution of the marginals of the –valued generalized marked empirical process Hereafter, Continuous Mapping Theorem, and Theorem 4.1 in Cuesta–Albertos, Fraiman and Ransford [12] will be considered, leading to a distribution free test, based on the critical values of the random projection of the infinite–dimensional centered Gaussian measure in induced by the autocovariance operator of –valued innovation process
3.1 Central Limit Theorem
In the following lemma, Theorem 2.16 in Bosq (2000) is formulated. We will apply it in the proof of Theorem 1 below.
Lemma 2
Let be an –valued martingale difference, and be an orthonormal basis of Suppose that
where is a family of real numbers, and
with Then,
where denotes the convergence in distribution. Here, the covariance operator of the functional Gaussian random variable satisfies
Without loss of generality, we will consider the special case of the orthonormal basis given by the eigenvectors of the autocovariance operator of
In the following result, the following aditional condition is also considered:
Assumption A3. Assume that and that (i.e., is in the trace class).
Theorem 1
Let be a centered ARH(1) process. Under Assumptions A1–A3, for each the functional empirical process (10) satisfies as Here, for every
In particular, considering for every
Proof.
The proof is based on the verification of the conditions assumed in the CLT for Hilbert–valued martingale difference sequences given in Theorem 2.16 in Bosq (2000) (see also Lemma 2 above). Let us first verify equation (11) in Lemma 2, corresponding to condition (2.59) in Theorem 2.16. For every and the following events are considered:
Clearly, and for every
From equation (15), applying Chebyshev inequality and stationarity, keeping in mind that the events and coincide for and we obtain
Dominated Convergence Theorem yields to
On the other hand, for any and applying SWN property of the innovation process under Assumptions A1–A2, from equation (5),
Thus, is uniformly integrable, and equation (11) is therefore satisfied.
Let us now prove that condition (2.61) in Theorem 2.16 in Bosq (2000) holds (see also equation (13) in Lemma 2). Specifically, considering again Assumptions A1–A2, from equation (5), and from the SWN property of applying Markov Theorem, for any
From (19), in view of the trace property of operator
Thus, equation (13) is satisfied.
Finally, we prove condition (2.60) in Theorem 2.16 in Bosq (2000) (equation (12) in Lemma 2) holds. Specifically, we prove that condition (2.36) in Corollary 2.3 in Bosq (2000) holds. Under Assumptions A1–A2, from equations (7) and (8), for every
Let us denote for Here, as before, for every
Under Assumptions A1–A3, from (5), applying several times Cauchy–Schwartz inequality for different inner and semi–inner products, as well as well known properties of conditional expectation, for each denoting by the orthonormal system of eigenfunctions of we obtain
Note that under Assumption A1, and, under Assumption A3, and Hence, we can apply Corollary 2.3 in Bosq (2000) with Thus, for all and each
leading to equation (12) in Lemma 2 as we wanted to prove. Thus, the weak convergence to the indicated –valued Gaussian random variable holds.
3.2 Functional Central Limit Theorem
Let be the separable Banach space of –valued continuous functions on with respect to the norm, under the supremum norm The following result provides an invariance principle for martingale difference sequences evaluated in a Hilbert space (see Theorem 2 in Walk [47]).
Theorem 2
Let be a martingale difference sequence of –valued random variables, with respect to the filtration satisfying Let be a trace operator. For each denote as the autocovariance operator of given That is,
Assume that
For
Then, the sequence of random elements in with the supremum norm, which are defined by
converges in distribution to a Brownian motion in with a.s., and covariance operator of
In the derivation of the next result, the following additional condition is assumed.
Assumption A4. where denotes here the sample functional value in corresponding to of innovation process at time for every
Theorem 3
For each let be the –valued martingale difference with respect to the filtration introduced in Lemma 1 (see also equation (23)), satisfying Assumptions A1–A4. Define, for each fixed the sequence in given by
where is the functional empirical process introduced in (10). Then, for each fixed as
with being an –valued Gaussian process, satisfying a.s., and with covariance operator of That is, is an –valued Gaussian process with independent and stationary –valued Gaussian increments for every In parti-cular, for every Thus, for each fixed a.s. lies in and it can be identified, in distribution sense, with –valued Brownian motion with covariance operator (see, e.g., Definition 2 in [15]).
Proof. The proof is based on verifying that, for each fixed the sequence in (23) satisfies Theorem 2(i)–(iii).
We first verify Theorem 2(i), under Assumption A1, keeping in mind (5), by computing for each and
Theorem 2(ii) holds, under Assumptions A1–A2, from equation (7), since, for each and every
Finally, Theorem 2(iii) also holds under Assumptions A1, A2 and A4. Specifically, we prove convergence in (see Remark 1). Thus, from equation (5), for each and applying Chebyshev’s inequality, we obtain as
since
Remark 2
From equation (9), in a similar way to equation (25), for every as
Thus, from Theorem 3, the –valued generalized Gaussian random field can be identified, in distribution sense, with –valued two–parameter Brownian Sheet with respect to function parameter (generalized random field), whose associated probability measure is and with respect to time parameter with associated probability measure given by the uniform distribution on the interval
One can look at the generalized marginal of the two–parameter Brownian sheet introduced, which can be identified in distribution sense with the –valued generalized Gaussian process providing the extension to the infinite–dimensional and generalized setting of the limiting Gaussian process (indexed by the co–variate), arising in Theorem 2.1 in [41], for the real–valued case under alternative conditions. In particular, from equations (9) and (27),
This process can then be identified in distribution sense with –valued generalized (indexed also by H) Gaussian Brownian motion, associated with probability measure on the parameter space (playing the role of measure in [41]). Equivalently,
where in the last identity, denotes transformed –valued Brownian motion. (see, e.g., Definition 2 in [15]).
From Theorem 3 and Remark 2, applying continuous mapping theorem, under the almost surely continuity, in the norm of of the sample paths of the Brownian Sheet with respect to time argument or parameter as well as with respect to function parameter (also in the norm of ), one can obtain, from reverse triangle inequality,
where has been introduced in (LABEL:fclt).
Let us now denote with From Theorem 3 (see equation (29)), one would reject the null hypothesis if
exceeds an appropriate critical value obtained from the boundary crossing probabilities of the two–parameter –valued generalized Brownian sheet over the unit interval and the Hilbert space Since these probabilities are not longer available, we will look for some alternative expressions of our test statistics, leading to known asymptotic critical values, applying, for instance, the random projection method (see Theorem 4.1 in Cuesta–Albertos, Fraiman and Ransford [12]).
From equation (29), looking at the marginals, we obtain
We restrict our attention to the weak convergence in (31) of Note that the limiting –valued generalized centered Gaussian process has covariance operator
for every It can then be identified in law with a –varying scaling of –valued Gaussian random variable i.e., with for Here, as before, the Gaussian measure in satisfies conditions (a)–(c) in Definition 2 of [15].
Under Assumption A1, implying the strictly stationarity of ARH(1) process, it is sufficient to assume satisfies Carleman condition, to apply Theorem 4.1 in Cuesta–Albertos, Fraiman and Ransford [12]. Note that a sufficient condition, ensuring Carleman condition holds, would be the existence of the moment generating functional of Thus, one would reject the null hypothesis if
exceeds an appropriate critical value of obtained by projection of the infinite–dimensional centered Gaussian measure into where has been generated from a non–degenerated centered Gaussian measure in Measure will be approximated in a strongly consistent way under (see [4], Chapter 4) in practice, by the Gaussian measure with autocovariance operator given by (see, e.g., equation (3.13) in Bosq [4]):
where is the operator completely specified in and
4 Consistency
In the results derived in this section the following additional condition is assumed:
Assumption A3. and
We derive sufficient conditions that will imply the consistency of goodness of–fit–tests based on for testing against the alternative where is a known operator. By it should be understood that the –measure of the set is positive, with Let for Assume that for every if and only if Let for every and consider
Under conditions of Theorem 1, Corollary 2.3 in Bosq [4] can again be applied, in a similar way to the derivation of equation (21) in the proof of Theorem 1, but considering the Hilbert space instead of the Hilbert space leading to, for every and
almost surely, where expectation is computed under the alternative (). Specifically, we apply Corollary 2.3 in Bosq [4] for computed from –valued random variables
That is, under Assumption A3, the following identities and inequalities are considered, from applying Cauchy–Schwartz inequality in and the definition of the bounded operator norm on keeping in mind that for every
The uniform almost surely convergence in in Equation (33), obtained from equation (34), can be equivalently expressed as
Hence, equation (35) together with Theorem 3, and the fact that if and only if yield the consistency of the test in an infinite–dimensional framework. Furthermore, one can also consider the testing based on projection into the random function obtained from a non–degenerated zero–mean Gaussian measure in (as given in equation (32)), equation (34) leads, in particular, to the consistency of the test
The study of the asymptotic power of the above test deserves attention, and will be discussed in a subsequent paper.
5 Misspecified
Theorem 3 is useful for testing the simple hypothesis However, in practice, the situation is a bit different, since is unknown, and we have to replace by a consistent estimator in the computation of the functional values of the empirical process in equation (10). Thus, we consider
In this situation, additional assumptions are considered (see Chapter 8 in Bosq [4]), to ensure suitable asymptotic properties, and, in particular, consistency of the projection estimator
where
Assumption A5. The following conditions are satisfied by –valued process :
is a standard ARH(1) process with a strong –white noise and
The eigenvalues of satisfying for every are strictly positive (i.e., ), and the eigenvectors are known.
for every
Under Assumption A5, is bounded satisfying:
Under Assumption A5, and some additional conditions, like the ones assumed in Lemma 8.1(3), and Theorems 8.5–8.6 in Bosq [4] (see also Lemma 3 below), the strong–consistency of
in the space of bounded linear operators on holds. Specifically, Lemma 3 below summarizes such conditions.
Lemma 3
The following assertions hold:
If for some
Under Assumption A5, if is a Hilbert–Schmidt operator, then,
(38) provided
(39) -
Under Assumption A5, considering is a Hilbert–Schmidt operator, and is bounded, then, for all there exists an integer such that implies
where and only depend on and the finite–dimensional probabilities of
When eigenvectors are unknown, we consider, for each and such that Thus, and respectively denote the system of eigenvalues and eigenvectors of the empirical autocovariance operator Then, for ensuring strong–consistency in the space of the projection estimator
Assumption A6 below will be considered. Here, and and for each and
Assumption A6. The following conditions are assumed on ARH(1) process
The eigenvalues of satisfy
where as before, for every
For every and a.s.
Under Assumption A6, and additional conditions, like the ones given in Theorem 8.7–8.8 in Bosq [4] (see also Lemma 4 below), the strong–consistency of the projection estimator holds in the space
In Lemma 4 below, we will use the following notation:
Lemma 4
Under Assumption A5(i) and Assumption A6(i)–(ii), if is a Hilbert–Schmidt operator, and for some
then,
(42) - (ii)
Under the conditions of Lemma 4 (see Corollary 8.3 in Bosq [4]), the following lemma also provides the consistency in of the empirical predictor of regarding the theoretical one for every
Lemma 5
Under conditions in Lemma 4(i),
Furthermore, under conditions in Lemma 4(ii),
Finally, if
Under conditions of Lemma 4(ii), the following result provides the asymptotic equivalence in probability in the norm of of and uniformly in where
Theorem 4
Under conditions of Lemma 4(ii), the following identity holds:
Proof.
Let us consider, for every
From Lemma 4(ii), Also, from equation (2.21) in Theorem 2.5(2) in [4], in a similar way to the proof of Theorem 3.9, see equations (3.36)–(3.37) in [4], Thus, (46) is obtained from (47).
Corollary 1
Under Assumptions A1–A6, and conditions of Lemma 4(ii),
From Corollary 1, one can proceed by testing in a misspecified framework in a similar way to equation (32) but in terms of Specifically, one would reject the null hypothesis if
exceeds an appropriate critical value of obtained by projection of the infinite–dimensional Gaussian measure into where, as in (32), has been generated from a non–degenerated centered Gaussian measure in
6 Final Comments
It is well-known that, from Theorem 1 in Cardot, Mas and Sarda [7], for the functional regression model with scalar response, and covariate taking values in a separable Hilbert space the projection estimator considered for the regression function does not satisfy a Central Limit Theorem (CLT) with convergence to a non–degenerate random element, in the topology induced by the norm of Thus, only weak–convergence results are obtained for the predictor, under suitable conditions, including suitable truncation order, and convexity of the eigenvalues of the autocovariance operator of the covariate at least at high frequency (in particular, their H2 condition on is crucial), among others. Then, a suitable normalization by the standard deviation of the partial sums of the sequence of independent centered random variables, involved in the prediction error, is considered. This Theorem 1 does not hold in the case of the ARH(1) model by the following two main reasons (see Mas [43]):
The converge to zero, in probability, of in the norm of the space of Hilbert–Schmidt operators on under suitable conditions like the ones assumed in Theorem 4.1, p. 98, in Bosq [4]. Here, denotes the projection operator into the subspace of generated by the eigenfunctions of In particular, under the conditions assumed in Theorem 4.1 in Bosq [4], our choice of the truncation order must be such that, as ensuring, in particular, Proposition 4 in Mas [43] holds. Note that Theorem 4.1 in Bosq [4] is proved by applying the strong law of large number to weakly dependent sequences of -valued random variables, that leads to the strong law of large numbers for ARH(1) processes (see, e.g., Theorem 3.7, p.86, in Bosq [4]). Specifically, from Lemma 4.1, p.96, in Bosq [4], on the AR representation of the diagonal self–tensorial product of an ARH(1) process, the strong consistency in the norm of of the empirical autocovariance operator of is obtained by applying Theorem 3.7 in Bosq [4].
The convergence in distribution of with to a centered Gaussian random Hilbert-Schmidt operator under the key condition This limit result is obtained by applying a CLT for an array of –valued martingale differences. The limit centered Gaussian random Hilbert-Schmidt operator has covariance operator defined by, for
(52) Note that, in the functional linear regression model with scalar response and functional covariate, we have no chance to formulate an equivalent condition to since the innovation process is real–valued. Hence, it is impossible to obtain convergence to a non–degenerate random element in the norm of (see Theorem 1 in [7]). A similar assertion holds for the functional linear regression model with functional response and covariate (see Theorem 8 in [11]). In this case, the independence of functional observations prevents from applying CLT for an array of –valued martingale differences.
Condition means that Thus, or, equivalently, belongs to the Reproducing Kernel Hilbert Space (RKHS) generated by the integral operator In the case where and have a common system of eigenfunctions, this condition can be equivalently expressed as
Hence, a faster decay of the eigenvalues of the autocovariance operator of the innovation process than the eigenvalues of the square autocovariance operator of the ARH(1) process is required. In particular, at least, the following identity holds
Section 2.3 in Mas [43] provides some examples of the eigenvalue sequence to define a suitable truncation order according to the assumed conditions for the derivation of his limit results above referred. In particular, one can consider if is of the form On the other hand, for one can choose if while if
In the case where the eigenfunctions of the autocovariance operator are unknown, projection into the empirical eigenfunctions of can be considered, and a CLT is obtained in Theorem 8.9 in Bosq [4]. Indeed, a similar decomposition to equation (11) in Cardot, Mas and Sarda [7] can be obtained. The strong consistency of the empirical eigenvalues and eigenvectors of is then applied (see Theorem 4.4, Lemma 4.3, Theorem 4.5, and Corollary 4.3 in Bosq [4]).
We recall that for the uniform in the norm of asymptotic equivalence in probability of the empirical processes and to hold in Theorem 4, the conditions of Lemma 4 must be considered, involving the quantities and in (LABEL:indisteigv). These quantities provide respective information about the decay velocity of the eigenvalues of and about the distance between such eigenvalues or their distribution. Both features of the pure point spectrum of must be taken into account in the model selection problem associated with a suitable truncation order In particular, if fast decay of the eigenvalues, like in example holds, and if the quantities in equation (LABEL:indisteigv), indicate a fast accumulation of the eigenvalues at a small neighborhood of zero, a slower divergence of as is required in order to ensure Theorem 4 is satisfied.
Finally, we highlight an interesting application of the approach presented to the context of goodness of fit tests for the family of spherical functional autoregressive processes, SPHAR(1) processes, which constitutes a special case of the SPHAR(p) class introduced in Caponera and Marinucci [6]. In particular, a consistent estimation of the isotropic autoregression kernel involved, and a Central Limit Theorem are derived in this paper. The extension of this approach to a special class of manifold functional autoregressive processes, in compact and connected two points homogeneous spaces, can also be achieved by applying the results in Ma and Malyarenko [42]. These two families of functional linear models open a new area of research where goodness of fit testing constitutes a challenging topic.
Ongoing work
Professor Manuel Febrero, co–author of this paper, is currently implementing a simulation study to illustrate the performance of the proposed goodness–of–fit test. This simulation study illustrates and complements the results derived in the Kolmogorov–Smirnov–type statistical testing framework, by numerical analyzing the finite sample performance of the corresponding continuous functional version of Cramér-–von Mises test statistics over the generalized (Hilbert–valued) empirical process via random projections.
Acknowledgements
This work has been supported in part by projects PID2022–142900NB-I00 financed by MCIU/ AEI/10.13039/501100011033 and by FEDER UE, and CEX2020-001105-M MCIN/AEI/10.13039/501100011033).