### v5

# Linear parametric model checking for functional time series

Abstract
The presented methodology for testing the goodness–of–fit of an Autoregressive Hilbertian model (ARH(1) model) provides an infinite–dimensional formulation of the approach proposed in [33], based on empirical process marked by residuals. Applying a central and functional central limit result for Hilbert–valued martingale difference sequences, the Wiener–type limiting process of the –valued empirical process indexed by an –valued covariate is obtained. The performance of the proposed testing procedure is illustrated by simulations. Consistency of this asymptotically distributed free test is derived. Since, in practice, the autocorrelation and autocovariance operators of the ARH(1) process are unknown, the uniform asymptotic equivalence in probability in –norm of the test statistics under totally and misspecified scenarios is established.
González–Manteiga, W., Ruiz–Medina. M.D., Febrero–Bande, M.
MSC2020. 60G10; 60G12; 60G18; 60G20; 60G22 (Primary). 60G60 (60G60).
Keywords. ARH(1) process, generalized functional empirical processes, goodness–of–fit test.
1 Introduction
Weakly dependent functional time series models have been extensively analyzed in the last few decades, supporting inference on stochastic processes, leading to the analysis of outstanding problems (see, e.g., [3]; [4]; [10]; [17]; [24]; [27]; [36]).
Hypothesis testing is still a challenging topic in the functional time series literature. Testing independence constitutes a central subject in this topic (see, e.g., [2]; [23]; [29]; [26]; [38]). On the other hand, second–order properties like separability, stationarity, and linear correlation range are usually tested to specify the covariance operator family (see, for instance, [9]; [28]; [31]; [32]; [39]). Alternative contributions in testing in functional time series literature, related, for example, with normality and periodicity tests can be respectively found in [21] and [25]. Two–sample mean and change point tests have also been formulated in a Banach space context to improve shape analysis for random elements (see, e.g., [15]).
In [26] a test is formulated, based on the sum of the -norms of the empirical correlation functions, to check if a set of a functional observations are independent and identically distributed. Consistency is analyzed as well. To detecting changes in the autocorrelation operator in functional time series, [27] derive a test based on change point analysis. In [28], several procedures are considered to test the null hypothesis of stationarity in the context of functional time series. The properties of the tests under several alternatives are also studied. Approaches for testing the structural stability of temporally dependent functional observations are considered in [40]. [31] propose a multistage testing procedure to determine the order of a functional autoregressive process. The test statistics involves the estimating kernel function in this linear model, and is approximately distributed as a chi–square with number of degrees of freedom determined by the number of functional principal components used to represent the data.
Bootstrap methods constitutes a challenging topic in functional time series to testing trend-stationary, under the umbrella of Kwiatkowski–Phillips–Schmidt–Shin (KPSS) tests (see, e.g., [7]). An interesting review on white noise tests in the context of functional time series is presented in [30], distinguishing between the formulation of such tests in the time or spectral domain, referring to the cases of strong white noise and general uncorrelated noise. The review also covers and extends several residual-based goodness-of-fit tests widely applied in functional data analysis. This survey provides, in addition, an interesting guide through simulations about implementation, benefits, and drawbacks of these tests.
In the context of functional regression special attention has been paid to significance testing (see, e.g., [5]; [8]; [22]). See also [20]. Some extensions to more complex hypothesis testing can be found in [1]; [14]; [19]; [35], among others. The last one adopts a marked empirical process based methodology, reducing the testing procedure to a multivariate problem by projection into the –sphere of A marked empirical process based approach is also adopted in [12] in the context of goodness of fit tests for the functional linear model with scalar response and functional covariate. The marked empirical process is indexed by random projections of the functional covariate. The weak convergence of the empirical process is obtained conditionally on a random direction. The almost surely equivalence of the test based on the original and projected covariate is proved. Calibration by wild bootstrap resampling is also achieved.
The present paper extends the formulation in [12] to the context of autoregressive functional time series, i.e., functional linear model with functional response, and time–dependent functional covariates, in the case of error term given by a Hilbertian Strong White Noise. Specifically, the following testing problem on the autocorrelation operator of an ARH(1) process is considered here:
where is a known operator. To address this problem, the following infinite–dimensional formulation of the involved marked empirical process indexed by a functional covariate is adopted, for each
where is an orthonormal basis of the separable Hilbert space and denotes the autoregressive process introduced in equation (1) below. The limit behavior of the marginals of process (1) under the null hypothesis is obtained by applying a Central Limit result for Hilbert–valued martingale difference sequences (see Theorem 2.16 in [4]). A functional central limit result is also derived by applying a special case of an invariance principle based on Robbins-Monro procedure (see Theorem 2 in [37]). A more general formulation can be obtained from the conditional central limit theorem in Hilbert spaces derived in [13].
Under the functional central limit result derived leads to the identification in law of the limiting process of marked empirical process (1). This limiting process is generalized Gaussian process, with index set a separable Hilbert space taking values also in Specifically, it coincides in law with Hilbert–valued Brownian motion on Its function parameter space is interpreted as the measurable space with being the Borel –algebra in and denoting the marginal of the random initial condition of ARH(1) process which under the assumed strictly stationarity characterizes the infinite–dimensional probability distribution of the marginals of ARH(1) process The asymptotic critical values of the goodness–of–fit test are then given by the boundary crossing probabilities of this –valued Brownian motion on Theorem 4.1 in [11] is applied to explicitly compute these asymptotic critical values by considering Gaussian random projections. Such projections are generated in our case from a Gaussian measure on with autocovariance operator given by the autocovariance operator of the innovation process of the ARH(1) model. The simulation study achieved, based on Fast Bootstrap, shows that an improvement of the performance of the test is obtained when increasing the functional sample size and the number of random projections. Consistency of the formulated test is proved. The case of misspecified autocorrelation and autocovariance operators of the ARH(1) process is also addressed, from previous results on strong–consistency of the considered projection autocorrelation operator estimator, and of the empirical eigenvalues and eigenfunctions of the autocovariance operator of ARH(1) process (see [4]). The uniform asymptotic equivalence in probability in –norm of the totally specified and misspecified functional empirical processes is proved. Thus, the limiting process in law in both cases coincides. Hence, one can proceed in the same way in the asymptotic testing procedure under the totally and misspecified ARH(1) model scenarios.
The outline of the paper is the following. Preliminaries are provided in Section 2. The second–order properties of the Hilbert–valued martingale difference sequence involved in the definition of the marked empirical process (1) are analyzed in Section 3. Section 4 derives a Central Limit result to characterize the asymptotic behavior of the marginals of the –valued empirical process. A Functional Central Limit Theorem for the considered Hilbert–valued marked empirical process indexed by infinite–dimensional covariates is obtained in Section 5, leading to the design of the goodness–of–fit test. The proposed goodness–of–fit procedure is illustrated by simulations in Section 5.1. Namely, in this section, the critical values of the test statistics are approximated fron Fast Bootstrap ( bootstrap replicates), and for different functional sample sizes and number of random projections (based on repetitions), the empirical test size and the empirical test power are computed. Sufficient conditions for the consistency of the proposed goodness–of–fit test are given in Section 6 as well. The issue of misspecified autocorrelation and autocovariance operators of the ARH(1) process is addressed in Section 7. Final discussion is presented in Section 8. The Supplementary Material provides the state of the art about asymptotic analysis, some auxiliary technical results and the proofs of main results of this paper.
2 Preliminaries
In the following we denote by the space of zero–mean –valued random variables on the basic probability space with for every Let be a zero–mean ARH(1) process on the basic probability space satisfying
and the conditions given in Chapter 3 in [4] for the existence of a unique stationary solution. Hence, denotes the autocovariance operator of and its cross–covariance operator.
The –valued innovation process is assumed to be –valued Strong White Noise (–SWN) (see Definition 3.1 in [4]). That is, is a sequence of independent and identically distributed zero–mean –valued random variables with
and functional variance that coincides with the trace norm of the autocovariance operator of We restrict our attention to ARH(1) processes satisfying the following condition:
Assumption A1. is a strictly stationary ARH(1) process with SWN innovation such that
Under model (1), keeping in mind Assumption A1, the following orthogonality condition
holds for with denoting the almost surely (a.s.) equality.
Let us consider a centered non–degenerated Gaussian probability measure on From Theorem 4.1 in [11], under Assumption A1, if satisfies Carleman condition, a class in of Borel sets of defines a separating class of if it contains a set of positive –measure. Thus, Lemma 1(d) in [16] can be reformulated in an infinite–dimensional framework by considering, for each and for a given orthonormal basis of
leading to the definition of the separating class of sets Hence, under Assumption A1, one can equivalently express the orthogonality condition (3) as
for a given orthonormal basis of It is straightforward that the sequence is an –valued martingale difference with respect to the filtration where for (see Lemma 1 in the Supplementary Material). Along the paper, from (1), the following identity will be applied several times (see equation (3.11) in [4]), for every
3 Second–order properties of –valued martingale difference sequence
For every in the computation of the second–order moments of –valued martingale difference sequence the following assumption will be considered.
Assumption A2. Assume that is independent of for all
Note that, for the strongly integrability of the marginals of the –valued martingale difference sequence allows the computation of the means and conditional means of the elements of this sequence from their weak counterparts (Section 1.3 in [4]).
Under Assumptions A1–A2, for any and and can be computed as follows:
where we have applied for all under Assumptions A1–A2, and denotes the trace norm of a nuclear operator on Here, is the infinite–dimensional marginal probability measure induced by and is the event for each Keeping in mind equation (5), the autocovariance operator of is given by
where In (2), we have applied the strictly stationarity condition of under Assumption A1, leading to, for every
In a similar way, the covariance operator can be computed for every
where the notation must be understood componentwise, by projections, as follows from equation (4). That is, the infinite–dimensional vector has components defined by the respective minimum values between the two projections of the two functions
4 Central Limit Theorem
Theorem 2.16 in [4] (see also Lemma 2 in the Supplementary Material) is applied in this section in the derivation of Theorem 1, providing the Gaussian limit distribution of the marginals of the –valued generalized marked empirical process in (1).
The following condition is assumed in the derivation of Theorem 1.
Assumption A3. Assume that and that (i.e., is in the trace class).
Without loss of generality, we will consider in the next results the special case of the orthonormal basis given by the eigenfunctions of the autocovariance operator of
Theorem 1
Let be a centered ARH(1) process. Under Assumptions A1–A3, for each the functional empirical process (1) satisfies as with denoting the convergence in distribution. Here, for every
In particular, considering for every
5 Functional Central Limit Theorem
This section provides the convergence in law of the functional generalized empirical process (1) to an -valued generalized (indexed by ) Gaussian process. This result follows from a functional central limit theorem, derived here by applying a special case of an invariance principle based on Robbins-Monro procedure (see Theorem 2 in [37]). Lemma 3 of the Supplementary Material provides this result. In a more general setting, the convergence to –valued generalized Brownian Sheet is also obtained from this result under suitable conditions. Hereafter, Continuous Mapping Theorem, and Theorem 4.1 in [11] will be considered, to formulate a distribution free test, based on the critical values of the random projections, computed from an infinite–dimensional centered Gaussian measure, whose autocovariance operator coincides with the one of innovation process
In the derivation of the next result, Lemma 3 in the Supplementary Material is applied to the martingale difference sequences
under the following additional condition
Assumption A4. where denotes here the sample functional value in corresponding to of innovation process at time for every
Let be the separable Banach space of –valued continuous functions on with respect to the norm, under the supremum norm The following Functional Central Limit result is formulated for the sequence in introduced in equation (2) below.
Theorem 2
For each let be the –valued martingale difference sequence with respect to the filtration introduced in equation (1). We consider that Assumptions A1–A4 hold. Define, for each fixed the sequence in given by
where is the functional empirical process introduced in (1). Then, for each fixed as
with being an –valued Gaussian process, satisfying a.s., and with covariance operator of That is, is an –valued Gaussian process with independent and stationary –valued Gaussian increments for every In parti-cular, for every Thus, for each fixed a.s. lies in and it can be identified, in distribution sense, with –valued Brownian motion with covariance operator (see, e.g., Definition 2 in [13]).
Remark 1
From equation (4), in a similar way to the proof of Theorem 2 in Section S3.2 of the Supplementary Material (see, in particular, verification in this proof of Lemma 3(i) in the Supplementary Material), for every as
Thus, from Theorem 2, the –valued generalized Gaussian random field can be identified, in distribution sense, with –valued two–parameter Brownian Sheet with respect to function parameter (generalized random field), whose associated probability measure is and with respect to time parameter with associated probability measure given by the uniform distribution on the interval
One can look at the generalized marginal of the two–parameter Brownian sheet introduced, which can be identified in distribution sense with the –valued generalized Gaussian process providing the extension to the infinite–dimensional and generalized setting of the limiting Gaussian process (indexed by the co–variate), arising in Theorem 2.1 in [33], for the real–valued case under alternative conditions. In particular, from equations (4) and (3),
This process can then be identified in distribution sense with –valued generalized (indexed also by H) Gaussian Brownian motion, associated with probability measure on the parameter space (playing the role of measure in [33]). Equivalently,
where in the last identity, denotes transformed –valued Brownian motion (see, e.g., Definition 2 in [13]).
From Theorem 2 and Remark 1, applying Continuous Mapping Theorem, under the almost surely continuity in the norm of of the sample paths of the Brownian Sheet with respect to time argument or parameter as well as with respect to function parameter (also in the norm of ), one can obtain, from reverse triangle inequality,
where has been introduced in (2).
Let us now denote with From Theorem 2 (see (5)), one would reject the null hypothesis if
exceeds an appropriate critical value obtained from the boundary crossing probabilities of the two–parameter –valued generalized Brownian sheet over the unit interval and the Hilbert space Since these probabilities are not longer available, we will look for some alternative expressions of our test statistics, leading to known asymptotic critical values, applying, for instance, the random projection method (see Theorem 4.1 in [11]).
From equation (5), looking at the marginals, we obtain
We restrict our attention to the weak convergence in (7) of Note that the limiting –valued generalized centered Gaussian process has covariance operator
for every It can then be identified in law with a –varying scaling of –valued Gaussian random variable i.e., with for Here, as before, the Gaussian measure in satisfies conditions (a)–(c) in Definition 2 of [13].
Under Assumption A1, implying the strictly stationarity of ARH(1) process, it is sufficient to assume satisfies Carleman condition, to apply Theorem 4.1 in [11]. Note that a sufficient condition, ensuring Carleman condition holds, would be the existence of the moment generating functional of Thus, one would reject the null hypothesis if
exceeds an appropriate critical value of obtained by projection of the infinite–dimensional random variable into where has been generated from a non–degenerated centered Gaussian measure in Measure will be approximated in a strongly consistent way under (see [4], Chapter 4) in practice, by the Gaussian measure with autocovariance operator given by (see, e.g., equation (3.13) in [4]):
where is the operator completely specified in and
5.1 Illustration
In this section, an illustration of the performance of the proposed goodness–of–fit procedure is given by simulation. We restrict our attention to the particular case of testing versus in the ARH(1) model (1). The empirical type I error and power, based on repetitions of the procedure implemented, for functional samples sizes are computed. The critical values of the test statistics are approximated from Fast Bootstrap based on bootstrap replicates. As given in Section 5, the random projection methodology is implemented in such computations (see also [12]).
Under a Gaussian scenario, the functional values of an ARH(1) process are generated with support in the interval evaluated at temporal nodes. A Gaussian zero–mean strong–white noise innovation process having integral autocovariance operator with exponential covariance kernel
is generated. A Gaussian random initial condition is also generated, independently of the innovation process has exponential autocovariance kernel having the same scale parameter values, as the innovation process Equation (1) is then recursively generated involving the numerical approximation of the integral
where, under the alternative hypothesis the kernel of the integral autocorrelation operator is given by for To remove dependence from the random initial condition, a burn-in period of observations is considered. That is, for each functional sample size the above–described recursive approximation of the values is computed removing the values
The random projection of the functional covariate and marks is obtained. The empirical process
is then expressed in terms of such random projections as
where and are independent realizations of Gaussian processes respectively having the same covariance kernel as processes and being those processes, and approximated by their truncated, at term Karhunen–Loéve expansions. Note that here we have considered two independent realizations of different Gaussian measures in for random projection, but, in our particular ARH(1) scenario, it is suffcient to consider for random projection a common realization of a Gaussian measure in (see equation (8)).
Fast Bootstrap is implemented from the equation
where are independent and identically distributed standard normal random variables. The computation of the –value is obtained for each projection from
where bootstrap replicates have been considered. The numbers of projections tested is To obtain only one –value the False Discovery Rate is computed (i.e., the expected proportion of false positives among the rejected hypotheses).
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.038 | 0.026 | 0.030 | 0.030 | 0.026 | 0.032 | 0.026 | |
| 0.034 | 0.020 | 0.024 | 0.024 | 0.020 | 0.026 | 0.024 | |
| 0.050 | 0.054 | 0.060 | 0.060 | 0.068 | 0.070 | 0.056 |
| 1 | 2 | 3 | 4 | 5 | 10 | 15 | |
| 0.340 | 0.394 | 0.414 | 0.456 | 0.456 | 0.458 | 0.466 | |
| 0.400 | 0.506 | 0.562 | 0.596 | 0.608 | 0.692 | 0.702 | |
| 0.526 | 0.692 | 0.786 | 0.832 | 0.862 | 0.960 | 0.978 |
In Tables 1 and 2, one can respectively find the –values and power approximations computed, based on repetitions of the implemented testing procedure, for functional sample sizes The Confidence Interval, based on repetitions, obtained for is Specifically, Table 1 shows the rejection probability when is true. One can observe in Table 1 that the computed estimates of the test size are below the true parameter value for the functional sample sizes While, for the functional sample size an improvement is observed in the computed estimates of for all number of random projections. For the functional sample sizes analyzed, no patterns are observed regarding computed estimates and the number of random projections. The empirical power of the test, for the same functional sample sizes and number of repetitions, are displayed in Table 2. One can observe the increasing patterns of the empirical power over the increasing and Thus, an improvement in the performance of the test is obtained when increasing the functional sample size and the number of random projections.
6 Consistency
In the results derived in this section the following additional condition is assumed:
Assumption A3. and
We derive sufficient conditions that will imply the consistency of goodness of–fit–tests based on for testing against the alternative where is a known operator. By it should be understood that the –measure of the set is positive, with Let for Assume that for every if and only if Let for every and consider
Under conditions of Theorem 1, Corollary 2.3 in [4] can again be applied, in a similar way to the proof of Theorem 1 in the Supplementary Material (specifically, we refer to the part of the proof where the verification of equation (1.2) of Lemma 2 is achieved). Note that, here, we are considering the Hilbert space instead of leading to
almost surely, for every and Here expectation is computed under the alternative (). Specifically, we apply Corollary 2.3 in [4] for computed from –valued random variables
That is, under Assumption A3, the following identities and inequalities are considered, from applying Cauchy–Schwartz inequality in and the definition of the bounded operator norm on keeping in mind that for every
The uniform almost surely convergence in in Equation (1), obtained from equation (2), can be equivalently expressed as
Hence, equation (3) together with Theorem 2, and the fact that if and only if yield the consistency of the test in an infinite–dimensional framework. Furthermore, one can also consider the testing based on projection into the random function obtained from a non–degenerated zero–mean Gaussian measure in (as given in equation (8)), equation (2) leads, in particular, to the consistency of the test
The study of the asymptotic power of the above test deserves attention, and will be discussed in a subsequent paper.
7 Misspecified
Theorem 2 is useful for testing the simple hypothesis However, in practice, the situation is a bit different, since is unknown, and we have to replace by a consistent estimator in the computation of the functional values of the empirical process in equation (1). Thus, we consider
In this situation, additional assumptions are considered (see Chapter 8 in [4]), to ensure suitable asymptotic properties, and, in particular, consistency of the projection estimator
where
Assumption A5. The following conditions are satisfied by –valued process :
is a standard ARH(1) process with a strong –white noise and
The eigenvalues of satisfying for every are strictly positive (i.e., ), and the eigenvectors are known.
for every
Under Assumption A5, is bounded satisfying:
Under Assumption A5, and some additional conditions, like the ones assumed in Lemma 8.1(3), and Theorems 8.5–8.6 in [4], the strong–consistency of
in the space of bounded linear operators on holds (see Lemma 4 in the Supplementary Material summarizing such conditions).
When eigenvectors are unknown, we consider, for each and such that Thus, and respectively denote the system of eigenvalues and eigenvectors of the empirical autocovariance operator Then, for ensuring strong–consistency in the space of the projection estimator
Assumption A6 below will be considered. Here, and and for each and
Assumption A6. The following conditions are assumed on ARH(1) process
The eigenvalues of satisfy
where as before, for every
For every and a.s.
Under Assumption A6, and additional conditions, like the ones given in Theorem 8.7–8.8 in [4], the strong–consistency of the projection estimator holds in the space Specifically, Lemmas 4–6 in the Supplementary Material provide the conditions required to ensure the strong–consistency of the functional parameter estimator of as well as of the associated plug–in predictor. These results allow the derivation, in the following theorem, of the asymptotic equivalence in probability in the norm of of the functional empirical process (1) under totally and misspecified The last one contemplating the case of unknown autocorrelation operator and autocovariance operator
Theorem 3
Under conditions of Lemma 5 in the Supplementary Material, the following identity holds:
Proof.
Let us consider, for every
From Lemma 5 in the Supplementary Material, Also, from equation (2.21) in Theorem 2.5(2) in [4], in a similar way to the proof of Theorem 3.9 (see, in particular, equations (3.36)–(3.37)) in [4], Thus, (5) implies (4).
Corollary 1
Under Assumptions A1–A6, and conditions of Lemma 5 in the Supplementary Material,
From Corollary 1, one can proceed by testing in a misspecified framework in a similar way to equation (8) but in terms of Specifically, one would reject the null hypothesis if
exceeds an appropriate critical value of obtained by projection of the infinite–dimensional Gaussian measure into where, as in (8), has been generated from a non–degenerated centered Gaussian measure in
8 Discussion
Up to our knowledge, this paper constitutes the first attempt to formulate the goodness–of–fit testing problem, based on the empirical process methodology, in the context of ARH(1) processes. The difficulties arising in the asymptotic analysis under misspecification of the autocorrelation and autocovariance operator of the ARH(1) process constitute one of the main reasons why this challenging topic has remained open (see Section S1 of the Supplementary Material). The present paper has solved these difficulties in the context of ARH(1) processes by applying the strong consistency results derived in [4] for projection estimators of the autocorrelation operator, including the case of unknown autocovariance operator. This last case was addressed in [4] by proving the strong consistency of the eigenvalues and eigenfunctions of the empirical autocovariance operator of an ARH(1) process under suitable conditions.
Finally, we highlight an interesting application of the approach presented to the context of goodness of fit tests for the family of spherical functional autoregressive processes, SPHAR(1) processes, which constitutes a special case of the SPHAR(p) class introduced in [6]. In particular, a consistent estimation of the isotropic autoregression kernel involved, and a Central Limit Theorem are derived in this paper. The extension of this approach to a special class of manifold functional autoregressive processes, in compact and connected two points homogeneous spaces, can also be achieved by applying the results in [34]. These two families of functional linear models open a new area of research where goodness of fit testing still constitutes a challenging topic (see also [18] for related asymptotics).
9 Supplementary Material
We have included in the Supplementary Material the state of the art on asymptotic analysis, some auxiliary technical results and the proofs of main results of this paper.
Acknowledgements
This work has been supported in part by projects PID2022–142900NB-I00 and PID2020-116587GB-I00, financed by MCIU/AEI/10.13039/501100011033 and by FEDER UE, and CEX2020-001105-M MCIN/AEI/10.13039/501100011033), as well as supported by grant ED431C 2021/24 (Grupos competitivos) financed by Xunta de Galicia through European Regional Development Funds (ERDF).